A car is 137 m from a stop sign and traveling toward the sign at 25 m/s. At this time, the driver suddenly realizes that she must stop the car. If it takes 0.200 s for the driver to apply the brakes, what must be the magnitude of the constant acceleration of the car after the brakes are applied so that the car will come to rest at the stop sign?
Answer #1

The distance travelled by the car in retardation motion is given by

s = 137-25*0.2= 137-5 = 132 m

The acceleration can be determined using

v^2-u^2 = -2as \\ a = u^2/2s = 25^2/264 = 2.37m/s^2